OpenStudy (anonymous):

How would you simplify 2 with an exponent of -3 into a fraction?

OpenStudy (mathmale):

The general rule to apply here is that a^(-1) = 1/(a) (the reciprocal of a). a^(-3) can be written as 1/(a^3). Try evaluating 2 with exponent of -3.

OpenStudy (precal):



OpenStudy (shamil98):

\[\large x^{-n} = \frac{ 1 }{ x^n }\] \[\large 2^{-3} = \frac{ 1 }{ 2^3 } = \frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

Thank you so much! That I understood best! :) @shamil98

OpenStudy (precal):

sorry so hard to use my mouse. I usually have better luck on my other computer